# Merging of Kolmogorov-Arnold networks trained on disjoint datasets

### v1

@@ 4.2  Scaling test
In the single-threaded setting for predicting determinants of $4\times 4$ matrices, approximately $10$ passes through the training dataset (epochs) are required to reach the Pearson correlation of at least $97\%$. This corresponds to processing $10^{6}$ records, and this value is used as the fixed workload for all parallel configurations. The number of threads is varied from $1$ to $6$ (the particular Intel CPU has $6$ performance ‘fast’ cores and $10$ efficiency ‘slow’ cores). Batch sizes range from $50\mathrm{K}$ down to $3\mathrm{K}$ records. For each combination of thread count and batch size, the required number of model merges (rounds) is chosen such that the total number of processed records remains exactly $1\mathrm{M}$.
@@
Under ideal scaling, the accuracy would remain unchanged; however, the merging of the models leads to a computational error, and the actual value drops from $96.8\%$ for $1$ thread to $94.5\%$ for $6$ threads. Such loss in accuracy can be compensated of course by increasing the number of rounds. For example, using $6$ threads, $20$ rounds, and the batch size of $16,666$ (twice larger computational work than in the table) yields the Pearson correlation of $97.0\%$, while completing in $2.22$ seconds — $2.4$ times faster than for $1$ thread.
@@
Table 2: Accuracies and execution times of the code on a laptop for the example of matrix determinants — strong scaling test.
4.3  Second example
The second example is not strictly a scaling test, as the computational workload varies, but it serves as a benchmark of parallel training to a fixed accuracy. The training is done on synthetic data, where the outputs are the areas of triangular faces of randomly-generated tetrahedra ($4$ faces), and the inputs are the coordinates of the vertices of the tetrahedra in 3D ($4$ vertices, $3$ coordinates for each). Number of the training records is $500\mathrm{K}$, number of the validation records is $50\mathrm{K}$. Three-layered KANs are trained, the first layer consisting of $70$ operators with $3$ points per function, the second layer consisting of $30$ operators with $18$ points per function, the third layer consisting of $4$ operators with $22$ points per function (i.e. the total of $42,960$ model parameters). MATLAB’s neural networks are used for the comparison with $4$ hidden layers and $80$ neurons per layer ($20,804$ model parameters). The results are shown in table 3, where it is seen that when trained to the same accuracy, parallel implementation of KANs is about $2.8$ times faster than the sequential and two orders of magnitude faster than MLP training on CPU.
Pearson (%) for outputs time (s)
MLP, CPU $98.6\pm 0.7$ $98.2\pm 0.9$ $98.9\pm 0.3$ $97.9\pm 0.4$ $1126.4\pm 69.5$
MLP, GPU $98.7\pm 0.5$ $98.3\pm 0.7$ $98.2\pm 0.8$ $98.3\pm 0.9$ $76.9\pm 1.2$
KAN sequential $8.6\pm 0.6 99$$9.0\pm 0.4$ $98.1\pm 0.2$ $98.5\pm 0.3$ $29.2\pm 3.2$
KAN parallel, 6 threads $98.5\pm 0.9$ $98.1\pm 0.7$ $98.6\pm 0.3$ $98.3\pm 0.4$ $10.5\pm 1.3$
Table 3: Baseline accuracy and performance tests for modelling of the areas of faces of random tetrahedrons conducted for MATLAB’s neural networks and the authors’ KANs’ implementation.
5  Experimental part performed on cluster — weak scaling
To assess the scalability of the code, the tests were run on an HPC cluster. All jobs were run on a single node equipped with two Intel(R) Xeon(R) Platinum 8358 CPUs ($32$ cores per CPU) and $256\mathrm{GB}$ RAM. The tests consisted in training KANs on datasets with the inputs being elements of $4\times 4$ or $5\times 5$ matrices and the outputs being the determinants of the matrices.
@@ 5  Experimental part performed on cluster — weak scaling
The results are shown in table 4, where it is seen that the scaling is very close to linear. For the example of $4\times 4$ matrix determinants, the efficiency is above $93\%$ when up to $32$ threads are used. For the example of $5\times 5$ matrix determinants, the efficiency is above $95\%$ when up to $16$ threads are used and only drops down to $83\%$ when $32$ threads are used (and even further for $64$ threads). Such noticeable decline in the latter case is due to a large number of model synchronisations and large model sizes for $5\times 5$ example.
In the example of $4\times 4$ matrix determinants, the datasets consisted of $100\mathrm{K}$ training records, the models consisted of $70$ addends, with $3$ points per inner function and $25$ points per outer function (i.e. the total of $5,110$ parameters per model). The batch size was taken to be $100\mathrm{K}$ records, i.e. the entire training dataset. The jobs consisted of $10$ batches, i.e. $10$ synchronisations (of $n$ models, where $n$ is the number of threads) are made within each execution of the code.
In the example of $5\times 5$ matrix determinants, the datasets consisted of $10\mathrm{M}$ training records, the models consisted of $160$ addends, with $3$ points per inner functions and $30$ points per outer function (i.e. the total of $16,800$ parameters per model). The batch size was taken to be $50\mathrm{K}$ records. The jobs consisted of $200$ batches.
@@
time (s), $4\times 4$ $6.64$ $6.80$ $6.82$ $6.82$ $7.11$ $7.04$ $7.79$
time (s), $5\times 5$ $217.02$ $220.24$ $226.07$ $222.82$ $227.68$ $260.99$ $282.96$
Table 4: Execution times of the code on HPC cluster for examples of matrix determinants — weak scaling test. The values are averages of three programme runs.
6  Concurrent-serial alternations of Newton-Kaczmarz
The conceptual description of the Newton-Kaczmarz algorithm provided in section 2 highlights its alternating concurrent-sequential structure. Each layer maps an input vector to an output vector through multiple independent operations that may be executed concurrently. However, the outputs of all concurrent computations must be synchronised before proceeding to the next layer. The backward updates follow the same pattern — a chain of potentially concurrent steps separated by unavoidable sequential dependencies.
These serial steps between independent computations represent a significant obstacle for concurrent training. Neither CPU nor GPU kernels are well suited for such patterns. This limitation can be illustrated using the classical vector-input-scalar-output two-layer KAN:
$\hat{z}=\sum_{j=1}^{n}A_{j},\qquad A_{j}=g_{j}\left(\sum_{l=1}^{p}f_{j,l}\left(x_{l}\right)\right),$ (8)
where $A_{j}$ stands for one addend, and each addend is itself a small tree mapping inputs $x_{l}$ to a scalar. All $n$ addends must be computed in order to obtain the residual $\Delta z$, and these computations can be executed concurrently. This residual, divided evenly among the addends and scaled by the learning rate, is the only quantity required for updating all functions $g_{j}$ and $f_{j,l}$. In the piecewise linear model, evaluating a function requires only two operations (one division and one multiplication), which is far too little computation for concurrency on CPUs or GPUs to provide any advantage. The synchronisation required after all $n$ addends are computed often takes longer than the time saved through parallelism.
However, this pattern is well suited for customised hardware, such as field-programmable gate arrays (FPGAs) or application-specific integrated circuits (ASICs). A CPU-based simulation demonstrating a custom synchronisation barrier designed by the authors is available online [^2] https://github.com/andrewpolar/detfast, but it serves only as a conceptual demonstration; synchronising at each training record introduces substantial overhead. The example implemented in the code is training a model for determinates of $5\times 5$ matrices. Training MLPs for the latter requires several hours [5], whereas the CPU-based KAN code completes it in a few minutes. An FPGA or ASIC implementation could reduce this further to seconds, because each addend can be mapped to a dedicated hardware processing block. In such customised hardware, concurrency is not emulated through threads or kernels but is realised through physical parallelism, where synchronization and data flow are governed directly by wiring rather than by software-controlled barriers.
6.1  Pretraining
Structure (8) motivates a pretraining strategy, in which groups of addends are trained concurrently over the entire dataset to predict an equal share of target $z$. For the example of determinants of $5\times 5$ matrices and a dataset of $10\mathrm{M}$ records, such pretraining produces a model with the Pearson correlation near $70\%$, providing a strong initialisation for the main training phase. When pretraining is combined with the concurrent-serial alternations, the full training process for $5\times 5$ determinants completes in just over $2$ minutes on the CPU used for the baseline examples.
@@ 7  Conclusions
The results demonstrate that KANs’ training can indeed be accelerated by iteratively training copies of the model on disjoint datasets (batches) and merging them. The speedup ratios between sequential and concurrent KAN training typically fall in the range of 3344, meaning that acceleration is possible, but there are some limits.
At the same time, even the sequential KAN implementation by the authors consistently outperforms MLPs on CPUs and, remarkably, also on a GPU. While direct CPU-GPU comparisons are formally inappropriate due to differences in hardware, these results are included for reference, taking into account that the GPU used for the benchmarking can be considered to be a typical laptop GPU (it has been noticed by the authors that some researchers have unrealistic expectations from GPU implementations).
Since methodology adopted in the authors’ code is FPGA-friendly, the authors believe that this direction is a perspective research capable to reduce hours into seconds of processing time for very large models and datasets.
MATLAB scripts [^3] http://openkan.org/downloadskan/matlabscripts.zip, sequential implementation [^4] https://github.com/andrewpolar/kankan-8, and parallel implementation [^5] https://github.com/andrewpolar/meanmodel are available online.

### v2

@@ 4.3  Scalability test 1: strong scaling, laptop CPU
In the single-threaded setting for Det4 example, approximately $10$ passes through the training dataset (epochs) are required to reach the Pearson correlation of at least $97\%$. This corresponds to processing $10^{6}$ records, and this value is used as the fixed workload for all parallel configurations. The number of threads is varied from $1$ to $6$ (the particular Intel CPU has $6$ performance ‘fast’ cores and $10$ efficiency ‘slow’ cores). Batch sizes range from $50\mathrm{K}$ down to $3\mathrm{K}$ records. For each combination of thread count and batch size, the required number of model merges (rounds) is chosen such that the total number of processed records remains exactly $1\mathrm{M}$.
@@
Under ideal scaling, the accuracy would remain unchanged; however, the merging of the models leads to a computational error, and the actual value drops from $96.8\%$ for $1$ thread to $94.5\%$ for $6$ threads. Such loss in accuracy can be compensated of course by increasing the number of rounds. For example, using $6$ threads, $20$ rounds, and the batch size of 16,66616{,}666 (twice larger computational work than in the table) yields the Pearson correlation of $97.0\%$, while completing in $2.22$ seconds — $2.4$ times faster than for $1$ thread.
Table 3: The strong scaling test for the authors’ C++ KANs’ implementation.
@@
4.4  Scaling test 2: weak scaling, HPC cluster
To assess the scalability of the code, the tests were run on an HPC cluster. All jobs were run on a single node equipped with two Intel(R) Xeon(R) Platinum 8358 CPUs ($32$ cores per CPU) and $256\mathrm{GB}$ RAM. Two examples are considered: Det4 and Det5.
@@ 4.4  Scaling test 2: weak scaling, HPC cluster
The results are shown in table 4, where it is seen that the scaling is very close to linear. The values are averages of three programme runs. For Det4 example, the efficiency is above $93\%$ when up to $32$ threads are used. For Det5 example, the efficiency is above $95\%$ when up to $16$ threads are used and only drops down to $83\%$ when $32$ threads are used (and even further for $64$ threads). Such noticeable decline in the latter case is due to a large number of model synchronisations and large model sizes for Det5 example.
In Det4 example, two-layer KANs are constructed, with $70$ blocks in the first layer and $3$ points per function, $1$ block in the second layer and $25$ points per function (the total of 5,1105{,}110 parameters per model). The batch size was taken to be $100\mathrm{K}$ records, i.e. the entire training dataset. The jobs consisted of $10$ batches, i.e. $10$ synchronisations (of $n$ models, where $n$ is the number of threads) are made within each execution of the code.
In Det5 example, two-layer KANs are constructed, with $160$ blocks in the first layer and $3$ points per functions, $1$ block in the second layer and $30$ points per function (the total of 16,80016{,}800 parameters per model). The batch size was taken to be $50\mathrm{K}$ records. The jobs consisted of $200$ batches.
Table 4: The weak scaling test for the authors’ C++ KANs’ implementation.
@@
time (s), Det4 $6.64$ $6.80$ $6.82$ $6.82$ $7.11$ $7.04$ $7.79$
time (s), Det5 $217.02$ $220.24$ $226.07$ $222.82$ $227.68$ $260.99$ $282.96$
4.5  Testing of the implementation on FPGA board
The implementation was run on Digilent Nexys A7-100T (Xilinx Artix-7) board. Two simple datasets are used: prediction of the Euclidean norms for random vectors of dimensions $4$ and $16$. The training dataset consists of $1024$ records, processed in $3$ cycles per record over $2$ epochs, for a total of $6144$ cycles. At a base clock frequency of $100\,\mathrm{MHz}$, the total execution time is $0.06144\,\mathrm{ms}$. The same number of latches is used for both the $4$-function and $16$-function models, as all forward-pass and update computations are performed concurrently.
@@ 5  Conclusions
The test results demonstrate strong performance consistency across MATLAB, FastKAN, and Keras, while the authors’ C++ KANs’ implementation delivers significantly faster training times. Even purely sequential implementation remains competitive with GPU-accelerated versions of the other software on a ‘standard’ laptop with CUDA support, and concurrent implementation achieves up to a $7$ times speedup.
Importantly, these benchmarks remain far from the theoretical performance limits. The computational flow of the considered KANs’ training method is exceptionally well-suited for hardware-supported concurrency available in FPGAs and ASICs, providing substantial potential for further acceleration. In such a hardware context, performance boundaries are determined not by software efficiency but by the size and capabilities of the hardware itself. Given ongoing industrial advances, KAN models trained using the proposed technique represent a highly promising avenue for scalable and high-performance AI.
Beyond raw training speed, maintainability and deployability are crucial considerations. The demonstrated C++ implementations are concise (300300700700 lines), avoid third-party libraries, and do not rely on language-specific features, making them straightforward to port to other programming environments and integrate into larger applications. This combination of performance, portability, and scalability positions such KANs’ training as a practical and impactful approach for future AI development.
MATLAB scripts [^1] http://openkan.org/downloadskan/matlabscripts.zip, a sequential C++ implementation [^2] https://github.com/andrewpolar/kankan-8, a parallel implementation [^3] https://github.com/andrewpolar/meanmodel, and register-transfer level (RTL) code [^4] https://github.com/andrewpolar/fpga for FPGAs are available online. The readers can also find useful discussions and simplified explanations of the presented concepts on the authors’ website [^5] http://openkan.org.
Acknowledgements
MP acknowledges funding from Research England’s “Expanding Excellence in England” (E3) fund via the “Multi-scale Multi-disciplinary Modelling for Impact” programme (M34Impact).